OpenStudy (anonymous):

Medal for this please Part A: Solve A = 9 over 2 (x + 23) for x. Part B: Determine the value of x when A = 108. Part C: Solve -np - 90 > 30 for n. Show your work

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

any idea?

OpenStudy (anonymous):

Part A: \[A = \frac{ 9 }{ 2(x+23) }\]

OpenStudy (anonymous):

Okay so first lets distribute the 2 to x and 23. What would you get?

OpenStudy (anonymous):

2x and 46

OpenStudy (anonymous):

is this for part B?

OpenStudy (anonymous):

No, this is for Part A. We have to solve for x. We haven't done that yet.

OpenStudy (anonymous):

Anyways so yes you get 2x + 46

OpenStudy (anonymous):

So now it looks like: \[A = \frac{ 9 }{ 2x + 46 }\]

OpenStudy (anonymous):

so do we add 46 and 9?

OpenStudy (anonymous):

No don't add.

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

so 414?

OpenStudy (anonymous):

then divide it by 2

OpenStudy (anonymous):

Wait hold up. Don't multiply either. I'll tell you.

OpenStudy (anonymous):

okay sorry haha

OpenStudy (anonymous):

Okay so lets multiply 2x + 46 on both sides. \[A(2x+46) = \frac{ 9 }{ 2x+46 }(2x+46)\]

OpenStudy (anonymous):

This cancels out 2x + 46 on the right side. And we have to then distribute the A to 2x and 46. \[2xA + 46A = 9\]

OpenStudy (anonymous):

So far so good?

OpenStudy (anonymous):

We have to find a common factor between 2xA and 46A. The common factor would be 2A since both coefficients are divisible by 2 and A. It would look like this now: \[2A(x+23) = 9\]

OpenStudy (anonymous):

We have to divide 2A on both sides. \[\frac{ 2A(x+23) }{ 2A } = \frac{ 9 }{ 2A }\]

OpenStudy (anonymous):

okay got it so far

OpenStudy (anonymous):

This cancels out 2A on the left side. Now the equation looks like this: \[x + 23 = \frac{ 9 }{ 2A }\] Now we can isolate x by subtracting 23 from both sides. \[x + 23 -23 = \frac{ 9 }{ 2A }-23\] Now, x solved is \[x = \frac{ 9 }{ 2A }-23\], which is the answer to PART A.

OpenStudy (anonymous):

For Part B, just substitute 108 for A in \[x = \frac{ 9 }{ 2A }-23\] What would you get for the value of x?

OpenStudy (anonymous):

its like doing part A but plugging in 108 instead of (x + 23)?

OpenStudy (anonymous):

Yes basically

OpenStudy (anonymous):

So what is the value of x for Part B?

OpenStudy (anonymous):

okay i can do it from here thank you :)

OpenStudy (anonymous):

Alright then. No problem:)